# Beyond Markovian: Reflective Exploration via Bayes-Adaptive RL for LLM Reasoning

### v1

@@ 4 The Necessity of Bayes-Adaptive RL for Reflective Reasoning
In a Bayes-Adaptive MDP (BAMDP) (bellman1959adaptive, ; martin1965some, ; lidayanbamdp, ), the agent maintains uncertainty over the underlying MDP, which is gradually reduced through interactions. Due to this implicit partial observability (duff2001monte, ; ghosh2021generalization, ), the policy and value depend on the full history $h_{t}$, instead of only the state $s_{t}$, to capture the agent’s evolving belief about the MDP parameters through cumulative observations. The objective for BAMDPs is
$\mathcal{J}_{\text{Bayes}}(\pi_{\theta}):=\mathbb{E}_{s_{0},\pi_{\theta}}\biggl{[}\sum_{t=0}^{T-1}\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid h_{t})}\bigl{[}r_{\mathcal{M}}(s_{t},a_{t})\bigr{]}\biggr{]},$
where $p(\mathcal{M}|h_{t})$ is the posterior distribution of $\mathcal{M}$ after observing $h_{t}$. This objective encourages the agent to not only maximize immediate rewards based on the current belief but also explore to gather more context about the uncertain MDP.
@@
Optimal adaptive policies naturally induce exploratory reflection behaviors, which provide additional contextual information even if the state remains identical. While reflective actions may be suboptimal relative to the (unknown) ground-truth MDP, the gathered context, especially the rewards, reduces the MDP’s uncertainty. This enables future policies to leverage the updated belief to act more optimally.
Comparison Between Markovian RL and Bayes-Adaptive RL Markovian RL depends on the current state, not history contexts, for decision-making. Exploration only happens during training with trial-and-error to discover return-maximizing action sequences, which are fully exploited as the optimal deterministic policy. In contrast, Bayes-Adaptive RL balances reward maximization with epistemic exploration throughout both training and testing, enabling meaningful test-time scaling with improved efficiency.
Theorem 4.2.
The test-time expected return of a Bayes-Adaptive policy can be exponentially higher in $T^{*}$ than that of the optimal Markovian policy, where $T^{*}$ is the minimal number of steps required to reach the correct answer under an optimal deterministic policy.
Sketch proof..
Consider the binary search tree in Figure 1. Let $s_{2}$, $s_{3}$, $s_{5}$, and $s_{6}$ be four equally possible candidate answers at test time, i.e., the prior $p(\mathcal{M}_{1})=p(\mathcal{M}_{2})=p(\mathcal{M}_{3})=p(\mathcal{M}_{4})=1/4$, where $r_{\mathcal{M}_{1}}(s_{2})=1$, $r_{\mathcal{M}_{2}}(s_{3})=1$, $r_{\mathcal{M}_{3}}(s_{5})=1$, $r_{\mathcal{M}_{4}}(s_{6})=1$, and all other rewards are zero. Here, $r(s)$ represents the reward of reaching $s$.
For any Markovian policy, the maximal return is $1/4$ (or $1/2^{d-1}$ for a depth-$d$ tree) since it is static and cannot adapt according to the reward feedback when reaching the four candidate answers. In contrast, the optimal Bayes-Adaptive policy has an expected return of $1$. This is achieved by updating the posterior $p(\mathcal{M}|h_{t})$ based on observation $h_{t}$ to eliminate the hypotheses $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\cdots$, until the ground-truth MDP has probability $1$, in which the agent can act optimally to reach the true answer. ∎
@@ 5 Method
The policy gradient for Bayes-Adaptive RL is as follows, which differs from (3.2) by replacing the value under a predefined $\mathcal{M}$ with a posterior-weighted value:
$\nabla_{\theta}\mathcal{J}_{\text{Bayes}}=\mathbb{E}_{s_{0},\pi_{\theta}}\biggl{[}\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid s_{t})\cdot\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid h_{t})}\bigl{[}Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\bigr{]}\biggr{]},$ (5.1)
where we write the history $h_{t}$ as $s_{t}$ in $\pi$ and $Q$ since in LLM reasoning, $s_{t}$ already encodes the full sequence of prior states and actions, and the reward is unavailable at test time so it’s absorbed into $\theta$ as a function of $s_{t}$. Here, we use the state-action value instead of the advantage since the latter requires multiple Monte Carlo rollouts at each step. By applying the Bayes rule, the posterior satisfies:
$p(\mathcal{M}\mid h_{t})=p(\mathcal{M}\mid s_{0:t},a_{0:t-1},r_{0:t-1})\propto p(\mathcal{M}\mid s_{0:t})\cdot p(r_{0:t-1}\mid s_{0:t},a_{0:t-1},\mathcal{M}),$ (5.2)
where $p(\mathcal{M}|s_{0:t})$ conditions only on the CoT, excluding rewards, and can be interpreted as the model’s probability of outputting solution $y_{s_{0}}^{\mathcal{M}}$. Interestingly, the second term $p(r_{0:t-1}|s_{0:t},a_{0:t-1},\mathcal{M})$ measures the likelihood of observing the rewards $r_{0:t-1}$ under $\mathcal{M}$ given the trajectory $s_{0:t}$. That is, we may write $p(r_{t}|s_{t},a_{t},\mathcal{M})\propto\exp(-\beta|r_{t}-r_{\mathcal{M}}(s_{t},a_{t})|)$ with a hyperparameter $\beta$ to obtain
@@
$\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid h_{t})}\bigl{[}Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\bigr{]}=\mathbb{E}_{\mathcal{M}\sim q(\mathcal{M}|s_{0})}\biggl{[}Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\frac{p(\mathcal{M}\mid h_{t})}{q(\mathcal{M}\mid s_{0})}\biggr{]}=\sum_{i=0}^{|\mathcal{M}|}Q^{\pi_{\theta}}_{\mathcal{M}_{i}}(s_{t},a_{t})p(\mathcal{M}_{i}\mid h_{t}),$ (5.4)
where the proposal $q(\mathcal{M}|s_{0})$ is the uniform distribution over the support of plausible MDPs, defined w.r.t. the ground-truth answer and candidate answers extracted from the model’s CoTs. We draw $|\mathcal{M}|$ CoT rollouts of $\pi_{\theta}$ on prompt $s_{0}$ to form $\{\mathcal{M}_{i}\}_{i=1}^{|\mathcal{M}|}$. The importance ratios are then self-normalized over $\{\mathcal{M}_{i}\}_{i=1}^{|\mathcal{M}|}$ and the ground-truth MDP $\mathcal{M}_{0}$ defined w.r.t. $y_{s_{0}}^{*}$. Substituting (5.2) and (5.3) into the above equation and letting $p(\mathcal{M}_{i}|s_{0:t})$ be the model’s state-conditional belief, we obtain:
$\small\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid h_{t})}\bigl{[}Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\bigr{]}=\sum_{i=0}^{|\mathcal{M}|}\underbrace{Q^{\pi_{\theta}}_{\mathcal{M}_{i}}(s_{t},a_{t})}_{\text{value in}\mathcal{M}_{i}}\cdot\underbrace{\pi_{\theta}(y_{s_{0}}^{\mathcal{M}_{i}}|s_{t}+\text{</think>})}_{\text{LLM's belief in }\mathcal{M}_{i}\text{'s plausibility}}\cdot\prod_{t^{\prime}=0}^{t-1}\underbrace{\exp\bigl{(}-\beta\bigl{|}r_{t^{\prime}}-r_{\mathcal{M}_{i}}(s_{t^{\prime}},a_{t^{\prime}})\bigr{|}\bigr{)}}_{\text{consistency b/w obs. \& }\mathcal{M}_{i}\text{'s pred.}},$ (5.5)
where $r_{t^{\prime}}$ is the actual observed reward as defined in (3.1), and $r_{\mathcal{M}_{i}}$, $Q_{\mathcal{M}_{i}}^{\pi_{\theta}}$ are defined in (3.3) w.r.t. the hypothesis MDP $\mathcal{M}_{i}$. For clarity, we have omitted the normalization constant when computing $p(\mathcal{M}_{i}|h_{t})$ from the last two terms. Algorithm 1 provides the unbatched version of the pseudocode.
@@
BARL offers a principled framework for stitching plausible strategies, analogous to linearized best-of-N reasoning, but with explicit step-level guidance on when and how LLMs should reflectively explore.
@@
6 How Bayes-Adaptive RL Helps Generalization: A Didactic Example
In this section, we present a didactic example to show how BARL facilitates test-time generalization. Consider the action space $\mathcal{A}$ that consists of three tokens, $\{0,1,2\}$, with one token generated at each timestep. The state is simply $s_{t+1}=a_{t}$. The objective is to repeat the prompt token three times consecutively within $29$ timesteps ($23^{3}+$ 2 is the minimal length of a sequence to include all unique triplets). The prompt token is 00 or $1$ at training time, and $2$ at test time. Episodes terminate when receiving a $1$ reward. This setup is illustrated in Figure 2.
@@ 6 How Bayes-Adaptive RL Helps Generalization: A Didactic Example
This synthetic task mirrors LLM reasoning, where the goal is not only to learn specific strategies (here, generating particular triplets) but also to acquire general problem-solving abilities, such as when and how to switch to new strategies. These capabilities are essential for handling distribution shifts between training and evaluation, a common challenge when developing effective reasoning models.
We use a $2$-head transformer encoder followed by a linear layer as the policy and train it using the policy gradient from Markovian RL (3.2) and from BARL (5.1). For Markovian RL, the value $Q^{\pi_{\theta}}$ in the policy gradient is $1$ only when $s_{0:T-1}$ contains the rewarding triplet $\mathop{\mathrm{argmax}}_{\text{tri}}r(\text{tri})$ of the ground-truth MDP, such as $000$ or $111$ during training. For BARL, we set $\beta=\infty$ so that $\prod_{t^{\prime}=0}^{t-1}\exp(-\beta|r_{t^{\prime}}-r_{\mathcal{M}_{i}}(s_{t^{\prime}},a_{t^{\prime}})|)=\mathbbm{1}(\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}_{i}}(\text{tri})\notin s_{0:t})$, i.e., the product is 00 when the rewarding triplet of $\mathcal{M}_{i}$ already appears in $s_{0:t}$ and thus is invalidated, [^2] This is because $r_{t^{\prime}}=0$ for unterminated sequences and if $\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}_{i}}(\text{tri})\in s_{0:t}$, i.e., $r_{\mathcal{M}_{i}}(s_{t^{\prime}},a_{t^{\prime}})=1$ for some $t^{\prime}$, then $\exp(-\beta|r_{t^{\prime}}-r_{\mathcal{M}_{i}}(s_{t^{\prime}},a_{t^{\prime}})|)=0$ and the resulting accumulated product is also 00. and $1$ otherwise. We let the policy’s state-conditional belief be $p(\mathcal{M}_{i}|s_{0:t})=1$ for $\mathcal{M}_{i}$ whose rewarding triplet aligns with the sampled policy action $a_{t}$, i.e., $\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}_{i}}(\text{tri})=s_{t-1:t+1}$, where $s_{t+1}=a_{t}\sim\pi_{\theta}(\cdot|s_{t})$. Therefore, the posterior-weighted value from (5.4) becomes
$\mathbb{E}_{\mathcal{M}}\bigl{[}Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\bigr{]}=\sum_{i=1}^{|\mathcal{M}|}Q^{\pi_{\theta}}_{\mathcal{M}_{i}}(s_{t},a_{t})p(\mathcal{M}_{i}\mid s_{0:t})\mathbbm{1}\bigl{(}\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}_{i}}(\text{tri})\notin s_{0:t}\bigr{)}=\mathbbm{1}\bigl{(}s_{t-1:t+1}\in\{\mathcal{M}_{i}\}_{i=1}^{|\mathcal{M}|},s_{t-1:t+1}\notin s_{0:t}\bigr{)},$
where $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$. The above formulation incentivizes the policy to eliminate hypotheses and switch to new strategies (i.e., new triplets) when the current strategy has been invalidated by earlier attempts up to step $t$, which is illustrated in Figure 3. This form of adaptive exploration provides a minimalist instantiation of BARL in synthetic settings. The difference between the above equation and (5.5) arises because the agent in this example is aware of the zero reward associated with unterminated episodes.
@@
We report the results in Figure 4, where accuracies are averaged over $50$ completions and the shadow regions are the standard deviation across $3$ independent model training runs. The results show that Markovian RL quickly finds and memorizes the training solutions but fails to generalize at test time. In contrast, Bayes-Adaptive RL increases both training and testing accuracies. Furthermore, its accuracy and convergence rate improve when given prior knowledge that rewarding triplets are repeated patterns, i.e., $|\mathcal{M}|=3$ with $r_{\mathcal{M}_{1}}(000)=r_{\mathcal{M}_{2}}(111)=r_{\mathcal{M}_{3}}(222)=1$ and all other rewards are zero. This highlights the advantage of more informative candidate sets, underscoring the importance of balancing the diversity and plausibility of the candidates. Specifically, they should be diverse enough to capture test-time uncertainty, yet constrained to only the most plausible candidates to shrink the hypothesis space.
Figure 4: (Left) Markovian RL (REINFORCE) memories training solutions and poorly generalizes. (Middle) BARL generalizes well at test time and (Right) improves with more informative candidate sets that strike a better balance between diversity and plausibility.

### v2

@@ 4 The Necessity of Bayesian RL for Reflective Exploration
In Bayesian RL, the agent maintains uncertainty over the underlying MDP, which is gradually reduced through interactions. Due to this epistemic partial observability (duff2001monte; ghosh2021generalization), the policy and value depend on the full history $h_{t}$, instead of only the state $s_{t}$, to capture the agent’s evolving belief about the MDP parameters through cumulative observations. The objective encourages the agent to not only maximize immediate rewards based on the current belief but also to explore to gather more context about the uncertain MDP.
Theorem 4.2 (Bayesian RL Admits Adaptive Optimal Policy).
For the Bayesian RL objective in (3.1) associated with any $(\mathcal{D},p(\mathcal{M}))$, there exists an optimal policy $\pi^{*}$ that is uncertainty-adaptive.
@@
The above result shows that optimizing the Bayesian RL objective in (3.1) attains an uncertainty-adaptive policy and thus naturally induces reflective exploration. While reflective actions may be suboptimal relative to the (unknown) ground-truth MDP, the gathered context, especially rewards, reduces the MDP uncertainty. Future policies can thus leverage the updated belief to act optimally.
Theorem 4.2 should be read as a structural result: for any $(\mathcal{D},p(\mathcal{M}))$, one can always represent a Bayes-optimal policy to be uncertainty-adaptive. It is possible that the dependence on the belief can be vacuous and the policy degenerates to Markovian. However, we will show in the following that there exist instances where any Bayes-optimal policy must be strictly uncertainty-adaptive, and the best Markovian policy can far underperform the optimal adaptive policy.
Theorem 4.3 (Gap Between Markovian and Adaptive Policies).
There exist instances $(\mathcal{D},p(\mathcal{M}))$ for which the objective in (3.1) is maximized by a strictly uncertainty-adaptive policy. Moreover, the performance gap between Markovian and strictly adaptive policies can be arbitrarily large.
Sketch of proof.
Consider the binary decision structure in Fig. 1. Let the prior $p(\mathcal{M}_{1})=p(\mathcal{M}_{2})=p(\mathcal{M}_{3})=p(\mathcal{M}_{4})=1/4$, where $r_{\mathcal{M}_{1}}(s_{2})=r_{\mathcal{M}_{2}}(s_{3})=r_{\mathcal{M}_{3}}(s_{5})=r_{\mathcal{M}_{4}}(s_{6})=1$ all other rewards are zero, and $\mathcal{D}=\varnothing$ is empty. Here, $r(s)$ represents the reward of reaching $s$. For any Markovian policy, the maximal return is $1/4$ (or $1/2^{d-1}$ for a depth-$d$ structure). In contrast, the optimal uncertainty-adaptive policy has a return of $1$. This is achieved by maintaining a hypothesis set $\mathcal{C}=\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}\}$. After visiting a leaf state and observing reward $r\in\{0,1\}$, update $\mathcal{C}\leftarrow\{\mathcal{M}\in\mathcal{C}:r_{\mathcal{M}}(s)=r\}$. Then each $r=0$ removes one hypothesis from $\mathcal{C}$, and $r=1$ identifies the true MDP. After at most four leaf visits, we have found the rewarding leaf and the return is $1$. We can increase $d$ to make the performance gap ($1$ versus $1/2^{d-1}$) arbitrarily large. ∎
Comparison Between Conventional RL and Bayesian RL Conventional RL does not admit strictly uncertainty-adaptive policies, which do not give rise to reflective exploration. Instead, exploration is trial-and-error during training and only serves to fit a fixed policy. In contrast, Bayesian RL yields uncertainty-adaptive policies whose actions depend on the evolving belief over MDPs, thereby incentivizing information gathering and inducing reflective behaviors through belief updates at deployment time.
Connection with Meta-RL and Continual Learning.
Although our Bayesian RL framework is orthogonal to meta-RL, we discuss the connections between them here. The goal of meta-RL is to learn a fast adaptation procedure across tasks, e.g., by optimizing deployment performance after only a few iterations of updates. From this perspective, uncertainty-adaptive policy in Definition 3.1 can be viewed as performing in-context learning (instead of parameter updates), and our framework can be interpreted as learning to do in-context exploration. Likewise, it is related to continual learning, in which the evolving belief over MDPs serves as an explicit, non-parametric memory that accumulates and reuses knowledge across tasks. However, our analysis focuses more on properties of policies that optimize the Bayesian RL objective, which enforce the Bayes-optimal solution to the exploration-exploitation trade-off. In other words, our framework explicitly maintains a belief over MDP and uses a belief-driven exploration strategy that rewards under the posterior.
@@ 5 Method
In the previous section, we analyzed the benefits of Bayesian RL and the role of self-reflection within the Bayesian RL framework. In what follows, we present a practical way to optimize the Bayesian RL objective in (3.1). The policy gradient for (3.1) is as follows, which differs from the conventional policy gradient by replacing the value under $\mathcal{M}^{*}$ with a posterior-weighted value:
$\nabla_{\theta}\mathcal{J}=\mathbb{E}_{s_{0},\pi_{\theta}}\biggl[\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}\mid h_{t})\cdot\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid\mathcal{D},h_{t})}\bigl[Q^{\pi_{\theta}}_{\mathcal{M}}(h_{t},a_{t})\bigr]\biggr].$ (5.1)
Here, we use the state-action value $Q$ instead of calculating the advantage since the latter requires branched Monte Carlo rollouts at each step, which is computationally costly. By applying the Bayes rule, the posterior satisfies:
$p(\mathcal{M}\mid\mathcal{D},h_{t})=p(\mathcal{M}\mid\mathcal{D},s_{0:t},a_{0:t-1},r_{0:t-1})\propto p(\mathcal{M}\mid\mathcal{D},s_{0:t})\cdot p(r_{0:t-1}\mid s_{0:t},a_{0:t-1},\mathcal{M}),$ (5.2)
where $p(\mathcal{M}|\mathcal{D},s_{0:t})$ conditions only on the CoT, excluding rewards, and can be modeled as the probability that the policy $\pi_{\theta}$, after training on $\mathcal{D}$, generates the solution $y_{s_{0}}^{\mathcal{M}}$. Interestingly, the second term $p(r_{0:t-1}|s_{0:t},a_{0:t-1},\mathcal{M})$ measures the likelihood of observing the rewards $r_{0:t-1}$ under $\mathcal{M}$ given the trajectory $s_{0:t}$. We follow ghosh2022offline to write $p(r_{t}|s_{t},a_{t},\mathcal{M})\propto\exp(-\beta|r_{t}-r_{\mathcal{M}}(s_{t},a_{t})|)$ with a hyperparameter $\beta$ to obtain
@@
$\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}\mid h_{t})}\bigl[Q^{\pi_{\theta}}_{\mathcal{M}}(h_{t},a_{t})\bigr]=\mathbb{E}_{\mathcal{M}\sim q(\mathcal{M}|s_{0})}\biggl[Q^{\pi_{\theta}}_{\mathcal{M}}(h_{t},a_{t})\frac{p(\mathcal{M}\mid h_{t})}{q(\mathcal{M}\mid s_{0})}\biggr]=\sum_{i=0}^{|\mathcal{M}|}Q^{\pi_{\theta}}_{\mathcal{M}_{i}}(h_{t},a_{t})p(\mathcal{M}_{i}\mid h_{t}),$
where the proposal $q(\mathcal{M}|s_{0})$ is the uniform distribution over the support of plausible MDPs, defined w.r.t. the ground-truth answer and candidate answers extracted from the model’s CoTs. We draw $|\mathcal{M}|$ CoT rollouts of $\pi_{\theta}$ on prompt $s_{0}$ to form $\{\mathcal{M}_{i}\}_{i=1}^{|\mathcal{M}|}$. The importance ratios are then self-normalized over $\{\mathcal{M}_{i}\}_{i=1}^{|\mathcal{M}|}$ and the ground-truth MDP $\mathcal{M}_{0}$ defined w.r.t. $y_{s_{0}}^{*}$. Substituting (5.2) and (5.3) into the above equation and letting $p(\mathcal{M}_{i}|s_{0:t})$ be the model’s state-conditional belief gives
[…]
where $r_{t^{\prime}}$ is the actual observed reward as defined in (3.2), and $r_{\mathcal{M}_{i}}$, $Q_{\mathcal{M}_{i}}^{\pi_{\theta}}$ are defined in (3.3) w.r.t. the hypothesis MDP $\mathcal{M}_{i}$. Since rewards are not available at deployment time, they are not explicitly provided as inputs to the policy. Instead, maximizing the Bayesian RL objective forces the parameters $\theta$ to encode a prediction of the reward as a function of state, thus absorbing the dependence on rewards into $\theta$. For clarity, we have omitted the normalization constant when computing $p(\mathcal{M}_{i}|h_{t})$ from the last two terms. Algorithm 1 provides the unbatched version of the pseudocode.
@@
BARL offers a principled framework for stitching together plausible strategies, analogous to linearizing best-of-N reasoning, but with guidance on when and how LLMs should reflectively explore.
@@
6 How BARL Helps Generalization: A Didactic Example
@@
Now we present a didactic example to show how BARL facilitates eest-time generalization. Consider the action space $\mathcal{A}$ that consists of three tokens, $\{0,1,2\}$, with one token generated at each timestep. The objective is to repeat the prompt token three times consecutively within $29$ timesteps ($3^{3}+2$ is the minimal length to include all unique triplets). The prompt token is 0 or $1$ during training, which does not have coverage for $2$. After training, the policy is deployed, where we consider the prompt token of $2$ as the evaluation task. Episodes terminate when receiving a $1$ reward. This setup is illustrated in Figure 2.
This setup mirrors LLM reasoning, where the goal is to not only learn specific strategies (here, generating particular triplets), but also general problem-solving abilities, such as when and how to switch to new strategies. These capabilities are essential for handling distribution shifts between training and eval, a common challenge when developing LLMs.
We use a $2$-head transformer encoder followed by a linear layer as the policy and train it using the policy gradient from conventional RL and from BARL (5.1). For RL, the state-action value $Q^{\pi_{\theta}}$ in the policy gradient is $1$ only when $s_{0:T-1}$ contains the rewarding triplet $\mathop{\mathrm{argmax}}_{\text{tri}}r(\text{tri})$ of the ground-truth MDP, such as $000$ or $111$ during training. For BARL, we set $\beta=\infty$ so that $\prod_{t^{\prime}=0}^{t-1}\exp(-\beta|r_{t^{\prime}}-r_{\mathcal{M}_{i}}(s_{t^{\prime}},a_{t^{\prime}})|)=\mathbbm{1}(\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}_{i}}(\text{tri})\notin s_{0:t})$, i.e., the product is 0 when the rewarding triplet of $\mathcal{M}_{i}$ already appears in $s_{0:t}$ (since $r_{t^{\prime}}=0$ for unterminated sequences) and $1$ otherwise. For the policy’s state-conditional belief $p(\mathcal{M}|s_{0:t})$, sampling from it is equivalent to sampling $a_{t}\sim\pi_{\theta}(\cdot|s_{t})$, with the associated $\mathcal{M}$ satisfying $\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}}(\text{tri})=s_{t-1:t+1}$. Therefore,
$\vskip 5.69046pt\mathbb{E}_{\mathcal{M}}\bigl[Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\bigr]=\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}|s_{0:t})}\bigl[Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})\mathbbm{1}(\mathop{\mathrm{argmax}}_{\text{tri}}r_{\mathcal{M}}(\text{tri})\notin s_{0:t})\bigr]=\mathbb{E}_{\pi}\bigl[\mathbbm{1}(s_{t-1:t+1}\notin s_{0:t})\bigr].$
The above formulation incentivizes the policy to eliminate hypotheses and switch to new strategies (i.e., new triplets) when the current strategy has been invalidated by earlier attempts up to step $t$, which is illustrated in Figure 3. This form of adaptive exploration provides a minimalist instantiation of BARL in synthetic settings. The difference between the above equation and (5.4) arises because the agent here is aware of the zero reward associated with unfinished episodes.
@@
We report the results in Figure 4, where accuracies are averaged over $50$ completions and the shadow regions are the standard deviation across $3$ independent model training runs. The results show that RL quickly finds and memorizes the training solutions but fails to generalize when deployed to the evaluation tasks. In contrast, Bayes-Adaptive RL significantly increases the evaluation accuracies. Furthermore, the accuracy and convergence rate improve when given prior knowledge that rewarding triplets are repeated patterns, i.e., $|\mathcal{M}|=3$ with $r_{\mathcal{M}_{1}}(000)=r_{\mathcal{M}_{2}}(111)=r_{\mathcal{M}_{3}}(222)=1$. This highlights the advantage of more informative candidate sets, underscoring the importance of balancing diversity and plausibility of the candidates. Specifically, they should be diverse enough to capture deployment-time uncertainty, yet constrained to only the most plausible candidates to shrink the hypothesis space.
Figure 4: Conventional RL (REINFORCE) memories training solutions and poorly generalizes beyond the training prompts. BARL performs well when deployed to the evaluation tasks and improves with more informative candidate sets.